OpenStudy (anonymous):

Nitrogen gas and hydrogen gas react to form ammonia (NH3) a) If the reaction mechanism is believed to 2 steps, propose a mechanism for the reaction if the intermediate species is H+. b) Write the net equation for this reaction. Make sure everything is balanced. c) Write the rate expression for this reaction

OpenStudy (anonymous):

so for (a) is it N2 + H2 = H+ ____ then H+____ = NH3

OpenStudy (anonymous):

This is called the haber process. By far the most important reaction ever realized. Provides energy for one third of the world. The Balanced equation is \[N_2 + 3H_2 = 2NH_3\] Therefore we can assume that if H+ is an intermediary it probably goes like this: \[N_2+3H_2 = 2N^+ + 6H^+ \]\[2N^++6H^+ = 2NH_3\] This, of course, is not the actual reaction mechanism since there is a platinum catalyst and the only time you can have ions as a gas like that are at extremely high temperatures but you asked for a two step solution.

OpenStudy (anonymous):

omg ty so would the net equation be N2+3H2=2NH3 and the rate expression= k(N2) (H2)^3 ?